OpenStudy (anonymous):

**WILL MEDAL** Which equations represent the line shown in the graph? Choose all answers that are correct. A.y+4=8/5(x+3) B.-8x+5y=17 C.y=8/5x+17/5 D.5x-8y=17

OpenStudy (anonymous):

a, c because the slope is 8/5

OpenStudy (anonymous):

Ok thx(:

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

Can u help me with one more plz?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Which equations in point-slope form are equations of the line that passes through the point (-7, 5) and (-4, 1)? CHOOSE EXACTLY TWO ANSWERS THAT ARE CORRECT. A.y-7=-3/4(x-5) B.y-5=-4/3(x+7) C.y+4=-3/4(x-1) D.y-1=-4/3(x+4)

OpenStudy (anonymous):

I think it's a and b @koook

OpenStudy (anonymous):

 since the slope is -4/3 the answer is b and d